A 100-kg man walks around a 500.0-m circular path in constant speed. Find:

 i. The radius of the circular path

 ii. The period of his motion if he completes 6 rounds in 30 minutes 

iii. The frequency of his motion

 iv. His speed

 v. The force that pulls him inward the path 




1
Expert's answer
2021-11-29T19:09:25-0500

i)

"r=L/(2\pi)=500/(2\pi)=79.62" m


ii)

"T=30\/6=5" min = "300" s


iii)

"f=1\/T=1\/300" Hz


iv)

"v=\omega r=2\pi fr=2\cdot 79.62\pi/300=1.67" m/s


v)

"F=ma=mv^2/r=100\cdot 1.67^2/79.62=3.49" N